OpenStudy (anonymous):

explain how to graph this linear equation. 4x+2y=4 @gypsy1274 2x-3y=10

OpenStudy (anonymous):

First solve each equation for y.

OpenStudy (anonymous):

i'll try my best 4x+2y=4 subtract (4x from both sides) 2y=4-4x (divide by 2) y=2-2x

OpenStudy (anonymous):

Yes. I would keep it in \(y=mx+b\) form to make graphing easier. \(y=-2x+2\) Start with the y-intercept. b=y-intecept

OpenStudy (anonymous):

please wait i was almost done and then messed up

OpenStudy (anonymous):

No problem. Take your time.

OpenStudy (anonymous):



OpenStudy (anonymous):

That looks right. The other equation is a little more difficult because you are dealing with fractions. I would still start by solving for y to determine the slope.

OpenStudy (anonymous):

the fractions are mostly what i need your help with.

OpenStudy (anonymous):

2x-3y=10 Solve for y:: What is your first step?

OpenStudy (anonymous):

2x-3y=10(subtract 2x from both sides -3y=10-2x y=10/ -3 -2x/-3

OpenStudy (anonymous):

Great. Can you determine the slope from that equation?

OpenStudy (anonymous):

slope =-3 and 1/3?

OpenStudy (anonymous):

\(2x-3y=10\) \(-3y=-2x+10\) \(y=\frac{2}{3}x+-\frac{10}{3}\) In \(y=mx+b\) form, it is easy to see what the slope should be. m = slope = \(\frac{2}{3}\)

OpenStudy (anonymous):

Does that make sense? Do you understand the \(y=mx+b\) form?

OpenStudy (anonymous):

y=2/3x+-10/3 so when you divide by negatives it has the same effect as multiplying?

OpenStudy (anonymous):

Yes. And even number of negatives makes a positive and an odd number of negatives makes a negative.

OpenStudy (anonymous):

For additional instruction on fractions, I would like to refer you to Purple Math http://www.purplemath.com/modules/fraction3.htm and Khan Academy www.khanacademy.org Both are free resources. And of course feel free to ask me questions at any time.

OpenStudy (anonymous):

im sorry can we continue using open study yahoo is too difficult for me to use

OpenStudy (anonymous):

OK, what ever works for you.

OpenStudy (anonymous):

ok im going to start a graph

OpenStudy (texaschic101):

y = 2/3x - 10/3 slope here is 2/3 y intercept is -10/3 or - 3 1/10 x intercept can be found by subbing in 0 for y y = 2/3x - 10/3 0 = 2/3x - 10/3 -2/3x = -10/3 x = (-10/3) / (-2/3) x = -10/3 * 3/2 x = -30/6 = -5 ---- x intercept is (-5,0) now plot your points...(-5,0) and (0,-10/3)...if it makes it easier to graph, the y intercept can be plotted as (0,-3 1/10)

OpenStudy (texaschic101):

oops...y intercept is -10/3 or -3 1/3 not -3 1/10..sorry bout that

OpenStudy (anonymous):